Shadow:

f(x) = 5x^2 - x^2 g(x) = 4x^2

Shadow:

\[\frac{ f(x) }{ g(x) }\] @satellite73

Shadow:

\[\frac{ 5x^2 - x^2 }{ 4x^2 }\]

satellite73:

you can cancel some stuff

satellite73:

are you sure there is not a typo there?

Shadow:

I'm trying to recall a problem from a quiz, so it's likely

satellite73:

\[5x^2-x^2=4x^2\] to start with

Shadow:

Oh got it, let me rewrite

Shadow:

f(x) = 5x^2 -x^1/3 g(x) = 4x^1/3

Shadow:

\[\frac{ 5x^2 - x ^{1/3} }{ 4x ^{1/3} }\]

satellite73:

you lost me

Shadow:

Basically wants me to simplify the above expression

satellite73:

you can rewrite it , but you can't "simplify" it

Shadow:

When you mean rewrite, you mean just \[- \sqrt[3]{x}\] For the top?

Shadow:

Let me just write a new question, haha sorry